# Coding Speech through Vocal Tract Kinematics

### v3

@@ VI Conclusion
We propose a novel encoding-decoding framework of speech, SPARC, which mimics the biophysical interface of speech production. Through large-scale training, our proposed SPARC framework achieves a high-performance articulatory inversion and synthesis that can generalize to unseen speakers, with a minimal loss of information compared to ground truth speech. To our knowledge, this is the first demonstration of universal articulatory synthesis that can scale up to an indefinite number of speakers. This universality is supported by a novel speaker identity encoding that embeds highly discriminable speaker-specific voice textures. Additionally, the encoded speaker embedding is effectively disentangled from articulatory features and can be used for accent-preserving voice conversion. Using SPARC, speech can be represented in a low-dimensional space, with each channel corresponding to an articulator in the vocal tract. This physical embodiment of speech allows for natural and intuitive control and interpretation.
@@ Acknowledgements
This research is supported by the following grants to PI Anumanchipalli — NSF award 2106928, BAIR Commons-Meta AI Research, the Rose Hills Innovator Program, UC Noyce Initiative, at UC Berkeley, and Google Research Scholar Award.
@@ A.1. Layer-wise AAI Performance
Since the EMA sensors are not spatially aligned across participants, each EMA channel from different participants is not directly comparable. As our model is using a particular participant’s (MNGU0’s) vocal tract as a template space, we find a mapping that spatially aligns MNGU0 sensors to different participants’ to calculate the correlation. Following [10], we pass speech audios from a target speaker from MOCHA-TIMIT or HPRC to our inversion model, and then a linear model is trained to predict the reference EMA of the target speaker. We used a linear regression with L1 penalty (Lasso) to impose sparsity in the weights, using $alpha=0.01$. Scikit-learn library is used for training the model (“sklearn.linear_model.Lasso”) [89]. This L1 penalty reveals channel-by-channel correspondence in the affine transformation (Fig. 9; more discussed in the next section). We split the data to 5 folds where 4 folds are used for training the linear model and 1 fold is used for evaluating the model. This is done individually for each speaker in MOCHA-TIMIT (7 speakers with 27 m of speech in average) and HPRC (7 speakers with 59 m of speech in average).
@@ A.5. Evidence for Excluding Coronal Axes of EMA
In our aritculatory encoding, we largely ignore the coronal axes of EMA data. Here, we provide evidence for such exclusion with a probing analysis. The same linear probing is conducted using SSL features (WavLM) and the original 12 midsagittal articulatory channels. We use EMA-MAE [90] for this experiment since the dataset includes two coronal channels: the left corner of the mouth (LC) and the lateral tongue (LT). Both for LC and LT, the linear probing using SSL features is significantly worse than using articulatory features (for LC, SSL features: $0.68\pm 0.19$, articulatory features: $0.73\pm 0.20$; for LT, SSL features: $0.67\pm 0.19$, articulatory features: $0.72\pm 0.17$). This suggests that those coronal features can be predicted with a fair correlation (0.72-0.73) by a linear combination of the existing 12 channels. These results suggest that the coronal axes can be safely ruled out for the proposed coding framework.

### v4

@@ VI Conclusion
We propose a novel encoding-decoding framework of speech, SPARC, which mimics the biophysical apparatus of speech production. Through large-scale training, our proposed SPARC framework achieves a high-performance articulatory inversion and synthesis that can generalize to unseen speakers, with a minimal loss of information compared to ground truth speech. To our knowledge, this is the first demonstration of universal articulatory synthesis that can scale up to an indefinite number of speakers. This universality is supported by a novel speaker identity encoding that embeds highly discriminable speaker-specific voice textures. Additionally, the encoded speaker embedding is effectively disentangled from articulatory features and allows accent-preserving voice conversion. Using SPARC, speech can be represented in a low-dimensional space, with each channel corresponding to an articulator in the vocal tract. This physical embodiment of speech allows for natural and intuitive control and interpretation.
@@ Acknowledgements
This research is supported by the following grants to PI Anumanchipalli — NSF award 2106928, BAIR Commons-Meta AI Research, the Hellman Fellows Program, the Rose Hills Innovator Program, UC Noyce Initiatives at UC Berkeley, and Google Research Scholar Award. We acknowledge Dr. Abdelrahman Mohamed (Rembrand) and Prof. Alan W Black (Carnegie Mellon University) for insightful discussions at various stages of this work.
@@ A.1. Layer-wise AAI Performance
Since the EMA sensors are not spatially aligned across speakers, each EMA channel from different speakers is not directly comparable. As our model is using a particular participant’s (MNGU0’s) vocal tract as a template space, we find a mapping that spatially aligns MNGU0 sensors to different participants’ to calculate the correlation. Following [10], we pass speech audios from a target speaker from MOCHA-TIMIT or HPRC to our inversion model, and then a linear model is trained to predict the reference EMA of the target speaker. We used a linear regression with L1 penalty (Lasso) to impose sparsity in the weights, using $alpha=0.01$. Scikit-learn library is used for training the model (“sklearn.linear_model.Lasso”) [89]. This L1 penalty reveals channel-by-channel correspondence in the affine transformation (Fig. 9; more discussed in the later section). We split the data to 5 folds where 4 folds are used for training the linear model and 1 fold is used for evaluating the model. This is done individually for each speaker in MOCHA-TIMIT (7 speakers with 27 m of speech in average) and HPRC (7 speakers with 59 m of speech in average).
@@ A.5. Evidence for Excluding Coronal Axes of EMA
In our aritculatory encoding, we largely ignore the coronal axes of EMA data. Here, we provide evidence for such exclusion with a probing analysis. The same linear probing is conducted using SSL features (WavLM) and the original 12 midsagittal articulatory channels. We use EMA-MAE [90] for this experiment since the dataset includes two coronal channels: the left corner of the mouth (LC) and the lateral tongue (LT). As shown in Fig. 10, both for LC and LT, the linear probing using SSL features is significantly worse than using articulatory features (for LC, SSL features: $0.68\pm 0.19$, articulatory features: $0.73\pm 0.20$; for LT, SSL features: $0.67\pm 0.19$, articulatory features: $0.72\pm 0.17$). This suggests that those coronal features can be predicted with a fair correlation (0.72-0.73) by a linear combination of the existing 12 channels. These results suggest that the coronal axes can be safely ruled out for the proposed coding framework.